# SignBot: Learning Human-to-Humanoid Sign Language Interaction

### v1

@@ Abstract
Sign language is a natural and visual form of language that uses movements and expressions to convey meaning, serving as a crucial means of communication for individuals who are deaf or hard-of-hearing (DHH). However, the number of people proficient in sign language remains limited, highlighting the need for technological advancements to bridge communication gaps and foster interactions with minorities. Based on recent advancements in embodied humanoid robots, we propose SignBot, a novel framework for human-robot sign language interaction. SignBot integrates a cerebellum-inspired motion control component and a cerebral-oriented module for comprehension and interaction. Specifically, SignBot consists of: 1) Motion Retargeting, which converts human sign language datasets into robot-compatible kinematics; 2) Motion Control, which leverages a learning-based paradigm to develop a robust humanoid control policy for tracking sign language gestures; and 3) Generative Interaction, which incorporates translator, responser, and generator of sign language, thereby enabling natural and effective communication between robots and humans. Simulation and real-world experimental results demonstrate that SignBot can effectively facilitate human-robot interaction and perform sign language motions with diverse robots and datasets. SignBot represents a significant advancement in automatic sign language interaction on embodied humanoid robot platforms, providing a promising solution to improve communication accessibility for the DHH community. Please refer to our webpage: https://qiaoguanren.github.io/SignBot-demo/
1 Introduction
@@
Sign language, as the primary linguistic medium for the deaf and hard-of-hearing (DHH) communities, plays a vital role in bridging their communication barriers with other people. Recent advancements in computer vision and large language models (LLMs) have significantly enhanced sign language applications, including generation, translation, and recognition [1; 2; 3; 4; 5]. These advancements enable effective translation between sign language text, videos, and mesh representations. However, despite these advancements, their real-world impact on assisting individuals with disabilities remains limited. A key reason is that these systems are primarily demonstrated in virtual environments and cannot facilitate physical interaction with people in real-world scenarios.
To bridge this gap, the recent development of Embodied Artificial Intelligence (EAI), which integrates AI models into a physical form, offers a promising solution. Unlike traditional sign language generation or comprehension models, EAI emphasizes real-world interaction, task execution, and continuous learning through physical experiences [6]. Recent advancements in EAI have extended to the development of humanoid robots [7; 8], which, due to their human-like mechanical structure, can seamlessly integrate into human workspaces and living environments. These robots are capable of performing various daily tasks, such as housekeeping [9; 10; 11], cooking [12], and navigation [13; 14], making them a valuable platform for fostering natural physical interactions with humans.
@@
Recent studies have explored learning-based perception and control algorithms, such as reinforcement learning (RL) [15; 16; 17] and imitation learning [18; 19], to develop control policies for humanoid robots. The commonly studied paradigms for learning-based methods include: 1) Teleoperation refers to the remote control of a humanoid robot by a human operator [11; 20; 21; 22; 23]. 2) Command-based Control means driving a humanoid robot with movement commands (e.g., velocity or affordance). Their control policies can be trained to imitate human behavioral norms [24; 16; 25; 26; 19]. However, no previous study has explored these methods for sign language applications. Teleoperation-based approaches typically rely on human manipulation, preventing robots from autonomously performing sign language, while command-based control methods primarily focus on the robot’s body without addressing the complexities of dexterous hand movements. Additionally, many dexterous robotic hands have limited degrees of freedom (DoFs), and the lack of wrist flexibility further restricts the accurate expression of the rich and diverse movements required for sign language.
@@
To overcome this challenge, we introduce SignBot, an expressive robotic sign language framework designed for seamless interaction with sign language users. Our framework integrates an embodied cerebellum, which controls the balance and movements of a humanoid robot, and a generative cerebral that is responsible for comprehending received sign language and generating corresponding gestures for communication. SignBot mainly consists of three key components: 1) Motion Retargeting, which maps the action sequences from human sign language datasets into a format compatible with robotic kinematics [16; 27]. 2) Policy Training (SignBot’s Cerebellum), which enables humanoid robots to first learn diverse sign language motions in a simulated environment with a decoupled policy [22; 23; 16]. Specifically, we utilize decoupled body policies to learn the entire sign language gesture. The upper body, including the hands, learns to track the target sign language poses through imitation learning, while the lower body maintains a stable standing posture using an RL policy. 3) Sign Language Interaction (SignBot’s Cerebral) integrates a sign language translator [28], a sign language responder [29], and a sign language generator [3], enabling the robot to understand user expressions and respond appropriately in sign language. To ensure stability while executing sign language gestures, we utilize a substantial amount of training data in the simulation to enhance its stability. By combining these elements, our framework enhances real-time human-robot interaction, bridging the communication gap between sign language users and embodied humanoid robot systems.
@@
We design various experiments to verify the performance of SignBot under Goal-Tracking and Cumulative Rewards metrics and visualize the poses of humanoid robots performing sign language. Experimental results show that SignBot exhibits accuracy, generalization, naturalness, and robustness with various datasets and robots. Overall, the main contributions of our paper are as follows:
@@
    Human-Robot Interaction for Minority. We propose an interactive sign language framework that enables seamless communication between humanoid robots and the DHH community.
@@
    Precise Sign Language Execution. Our humanoid robot control policy robustly adapts to a diverse range of human sign language motions, ensuring stable and accurate execution.
@@
    Embodiment and Domain Adaptation. Our method can be transferred to different robots, achieving Sim-to-Real deployment and applying robot sign language in real interaction scenarios.
2 Related Work
@@
Humanoid Robotic Imitation of Human Behavior. For the imitating human behavior problem of humanoid robots, researchers often adopt a whole-body control learning paradigm [7]. This paradigm consists mainly of two approaches. One approach is to decouple the upper and lower body policies, which are responsible for managing different parts of a humanoid robot. Representative works include Exbody [16], Mobile-Television [30], WoCoCo [31], OmniH2O [22], etc. Although upper and lower body policies are decoupled, they can still be integrated into a whole-body control paradigm. The alternative approach involves providing reference motions for the humanoid robot. Given the physical similarities to humans, a promising reference is the collection of human movements from motion datasets, such as H2O [20], Exbody2 [24], ASAP [25], UH-1 [32], [33], HiLo [34], HWC-Loco [35], etc. We simultaneously leverage the advantages of both methods to perform sign language. These reference motions provide rich signals for humanoid robots to imitate human-like motions.
@@
Sign Language Processing. The field encompasses two primary research directions: sign language translation (SLT), and sign language generation (SLG). SLT and SLG form complementary pathways for bidirectional communication between deaf and hearing populations, specializing in sign-to-text and text-to-sign conversion, respectively [3]. Some studies successfully incorporated language models (LMs) pre-trained on extensive natural language corpora into SLT frameworks, yielding substantial accuracy enhancements [1; 5]. Recent state-of-the art SLG works can be categorized into two classes: the first group of methods [36] employ diffusion models to generate sign motions conditioned on text inputs; the second group of methods [37; 4; 3] consider the linguistic nature of sign languages and adopts a tokenizer-LM autoregressive generation approach.
3 Problem Definition
@@
Robot Learning Environment. We formulate the task of tracking human sign language motions as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $\mathcal{M}=(\mathcal{O},\mathcal{S},\mathcal{A},P_{\mathcal{T}},\mathcal{R},\mu_{0},\gamma)$, where: 1) Within the observation space $\mathcal{O}$, each observation $o_{t}\in\mathcal{O}$ consists of two components: proprioception ($o_{t}^{p}$) and goal imitation ($o_{t}^{y}$). The proprioception $o_{t}^{p}$ includes essential motion-related information such as the root state, joint positions, and joint velocities. Meanwhile, the goal imitation $o_{t}^{y}$ represents a unified encoding of the whole-body sign language pose that must be tracked during reinforcement learning (RL) training. 2) $s_{t}\in\mathcal{S}$ records the complete information and environment of the robot. We summarize a state as $s_{t}=[o_{t},o_{t-1},...,o_{t-H}]$ and each $o_{t}=[o_{t}^{p},o_{t}^{y}]$. 2) $a_{t}\in\mathcal{A}$ denotes the action space, and action $a\in\mathcal{A}$ denotes the target joint positions that a PD controller uses to actuate the DoF. 3) $r_{t}=\mathcal{R}(s_{t},a_{t})$ denotes the reward functions, which typically consist of penalty, regularization, and task rewards. These reward signals determine the level of optimality in the control policy, for which we provide a detailed introduction in Section 4.2. 4) $P_{\mathcal{T}}\in\Delta^{\mathcal{S}}_{\mathcal{S}\times\mathcal{A}}$ denotes the transition function as a mapping from state-action pairs to a distribution of future states. 5) $\mu_{0}\in\Delta^{\mathcal{S}}$ denotes the initial state distribution. 6) $\gamma\in(0,1]$ denotes the discounting factor. Under this POMDP, our goal is learning a control policy $\pi:\mathcal{S}\to\mathcal{A}$ that can maximize the discounted cumulative rewards $\sum_{t=0}^{T-1}\gamma^{t}r_{t}$. During the training process, we apply the PPO-based [38] algorithm to optimize the policy parameters: $\pi:\mathcal{S}\to\mathcal{A}$ where $H$ is the length of history windows, intending to maximize the expected return of the discounted episode reward $J(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\gamma^{t}r_{t}\right]$, where $T$ represents the length of the horizon of an episode. We learn the control policy via KL-regularized PPO [39; 40], and the optimized objective function can be defined as:
$\mathbb{E}_{t}\left[\frac{\pi(a_{t}\mid s_{t})}{\pi_{\mathrm{old}}(a_{t}\mid s_{t})}\hat{A}(s_{t},a_{t})+\alpha\mathcal{H}(\pi(s_{t}))-\eta\mathrm{KL}(\pi_{\mathrm{old}}(s_{t}),\pi(s_{t}))\right]$ (1)
where $\hat{A}(s_{t},a_{t})$ denotes the estimated advantage function, $\mathcal{H}(\pi(s_{t}))$ denotes the entropy of policy and $\alpha,\eta$ are the regularization coefficients.
@@
Figure 1: Overview of SignBot: The framework consists of three stages: (1) Motion Retargeting aligns human sign language gestures with the body structure of humanoid robots (Section 4.1). (2) Cerebellum performs Sim2Real policy training that enables the robot to track various sign language gestures in the simulated environment and deploy the policy to real-world (Section 4.2). (3) Cerebral conducts sign language reasoning to facilitate communication with sign language users through the sign language translator, response, and generator within the cerebral (Section 4.3).
4 SignBot
In this section, we introduce the pipeline of SignBot, which is divided into three parts: 1) Motion Retargeting of the body and hands, 2) Policy Training to control the robot’s movements as "SignBot’s cerebellum", and 3) Sign Language Reasoning for comprehensive and responding users’ sign languages as "SignBot’s cerebral" [41]. Figure 1 presents the pipeline of SignBot.
4.1 Motion Retargeting
@@
Body Retargeting. Our humanoid body retargeting system is based on an improved version of [42]. By establishing a mapping relationship between the source keypoints and the target keypoints, we design a dual T-Pose (the standard poses of the source and target skeletons) as a spatial alignment reference. We convert the local quaternion of each joint to the form of axis angles. Then the mapped angle is presented: $\textbf{m}=\beta\textbf{a}$, where $\textbf{m}=[q_{i},q_{j},q_{k}]$ corresponds to the degrees of freedom in local quaternion, $\beta$ denotes the rotation angle, and $\textbf{a}=[a_{x},a_{y},a_{z}]$ is the rotation axis. For 1D joints (the elbow, torso, knee, and ankle), we take the rotation angle projected onto the corresponding rotation axis of the 1D joints. It is important to note that to make the robot’s sign language movements more natural, we add two additional degrees of freedom at the robot’s wrist and transform the wrists and shoulders represented by a 1D joint to a 3D joint. As shown in Figure 1, the retargeted humanoid motion data retains critical sign language characteristics, including gesture accuracy and expressiveness under our hardware system.
@@
Hand Retargeting. To map human hand pose data to the joint positions of the Linker hand (the robot hands used by SignBot), we apply the preprocessing method from [27], adapting it specifically for the Linker hand. This process is often formulated as an optimization problem [43], where the difference between the keypoint vectors of the human hand model and the dexterous hand is minimized. Due to the misaligned DoFs in the fingers of the linker hand with the human hand model, we properly freeze several finger joints. Since the linker hand is larger than a typical dexterous hand, we adjust the scale factor in the optimization process. Additionally, we modify the regularization term to smooth the sign language movements between consecutive frames.
4.2 SignBot’s Cerebellar: Control Policy Training
@@
Within our SignBot humanoid agent, the cerebellum controls the low-level movements for performing sign language. As shown in the second stage of Figure 1, we train a motion control policy to enable the humanoid robot to track and imitate sign language gestures in a simulated environment. In this section, we discuss our approach from three key perspectives: observation space, decoupled policy, and reward design.
@@
Observation Space. SignBot’s observation $o_{t}\in\mathcal{O}$ consists of proprioception ($o_{t}^{p}$) and goal imitation ($o_{t}^{y}$) (Section 3). Our proprioception is defined as $o_{t}^{p}\triangleq\left[\mathbf{q}_{t},\dot{\mathbf{q}_{t}},\mathbf{v}_{t},\mathbf{w}_{t},\mathbf{g}_{t},\mathbf{a}_{t-1}\right]$, which includes joint position $\mathbf{q}_{t}\in\mathbb{R}^{55}$ (DoF position), joint velocity $\dot{\mathbf{q}}_{t}\in\mathbb{R}^{55}$ (DoF velocity), root linear velocity $\mathbf{v}_{t}\in\mathbb{R}^{3}$, root angular velocity $\mathbf{w}_{t}\in\mathbb{R}^{3}$, root projected gravity $\mathbf{g}_{t}\in\mathbb{R}^{3}$, and the previous action $\mathbf{a}_{t-1}\in\mathbb{R}^{55}$. The goal observation is $\mathbf{o_{t}^{y}}\triangleq\left[\hat{\mathbf{q}}^{kp},\hat{\mathbf{q}},\dot{\hat{\mathbf{q}}}\right]$, where $\hat{\mathbf{q}}^{kp}\in\mathbb{R}^{14\times 3}$ are the positions of 14 selected reference keypoints [16] to ensure that the humanoid robot and the imitation goal are oriented in the same direction, $\hat{\mathbf{q}}\in\mathbb{R}^{55\times 3}$ are the positions of all reference joints, and $\dot{\hat{\mathbf{q}}}$ is the linear velocity of the reference joints. At the same time, we generally perform domain randomization [22; 20] in the simulation environment to ensure robustness and generalization. Appendix B.3 summarizes the domain randomization terms that we have utilized in SignBot.
@@
Decoupled Policy. Given that sign language involves precise coordination of both hand poses and full-body motion with high DoFs, learning a unified control policy is inherently challenging. More importantly, while the dual arms in the upper body can often be governed by a shared control strategy, the control approach for the lower body may vary significantly depending on the physical design of the humanoid robot. For instance, bipedal humanoid robots typically utilize reinforcement learning (RL) controller, whereas wheeled robots often employ model predictive control (MPC) techniques.
@@
Motivated by recent advances in whole-body humanoid control [16; 30], we adopt a decoupled architecture that separates the control policies of the upper and lower body.The primary objective of the upper-body policy $\pi^{upper}$ is to track the retargeted actions, while the lower-body policy $\pi^{lower}$ ensures balance in the robot’s default standing pose while adapting to the movements of $\pi^{upper}$.
@@
Where: 1) $a_{*}^{up}$ represents the retargeted action of the upper body, and $\delta$ denotes the Dirac delta function. 2) $\mathcal{D}_{f}(\cdot\|\cdot)$ refers to the $f$-divergence (e.g., KL-divergence) between two distributions. 3) The whole-body humanoid action, $a=[a^{low},a^{up}]$, is the concatenation of upper and lower body actions. 4) The reward from the humanoid control learning environment can be represented by the weighted penalty $r_{\mathfrak{P}}$, task $r_{\mathfrak{T}}$, and regularization $r_{\mathfrak{R}}$ terms [23]: $r_{t}=\beta_{\mathfrak{T}}r_{\mathfrak{T}}+\beta_{\mathfrak{P}}r_{\mathfrak{P}}+\beta_{\mathfrak{R}}r_{\mathfrak{R}}$. In our task, task rewards ($r_{\mathfrak{T}}$) measure the robot’s performance in tracking joint motions and body velocities. Penalty rewards ($r_{\mathfrak{P}}$) serve to discourage undesirable outcomes such as falling and violating dynamic constraints like joint or torque limits. Regularization rewards ($r_{\mathfrak{R}}$) are used to align the humanoid’s sign language gestures with human preferences. Appendix B sepcifies the reward functions.
@@
Note that to train the lower-body policy $\pi^{lower}$, we apply the PPO [38] algorithm for leggy humanoid robots and MPC for wheeled humanoid robots. In this way, SignBot can be scaled to multiple embodiments, for which we demonstrate in the experiment.
4.3 SignBot’s Cerebral: Sign Language Reasoning

### v3

@@ Abstract
Sign language is a natural and visual form of language that uses movements and expressions to convey meaning, serving as a crucial means of communication for individuals who are deaf or hard-of-hearing (DHH). However, the number of people proficient in sign language remains limited, highlighting the need for technological advancements to bridge communication gaps and foster interactions with minorities. Based on recent advancements in embodied humanoid robots, we propose SignBot, a novel framework for human-robot sign language interaction. SignBot integrates a cerebellum-inspired motion control component and a cerebral-oriented module for comprehension and interaction. Specifically, SignBot consists of: 1) Motion Retargeting, which converts human sign language datasets into robot-compatible kinematics; 2) Motion Control, which leverages a learning-based paradigm to develop a robust humanoid control policy for tracking sign language gestures; and 3) Generative Interaction, which incorporates translator, responder, and generator of sign language, thereby enabling natural and effective communication between robots and humans. Simulation and real-world experimental results demonstrate that SignBot can effectively facilitate human-robot interaction and perform sign language motions with diverse robots and datasets. SignBot represents a significant advancement in automatic sign language interaction on embodied humanoid robot platforms, providing a promising solution to improve communication accessibility for the DHH community. Please refer to our webpage: https://qiaoguanren.github.io/SignBot-demo/
I INTRODUCTION
@@
Sign language, as the primary linguistic medium for the deaf and hard-of-hearing (DHH) communities, plays a vital role in bridging the communication barriers between these communities and others. Recent advancements in computer vision and large language models (LLMs) have significantly enhanced sign language applications, including generation, translation, and recognition [1, 2]. These advancements enable effective translation between sign language text, videos, and mesh representations. However, despite these advancements, their real-world impact on assisting individuals with disabilities remains limited. A key reason is that these systems are primarily demonstrated in models and cannot facilitate physical interaction with people in real-world scenarios.
To address this gap, Embodied Artificial Intelligence (EAI) integrates AI models into physical agents, enabling real-world interaction, task execution, and continuous learning. Recent progress in humanoid robots [3] highlights their potential, as human-like structures allow seamless integration into daily environments for tasks such as housekeeping, cooking, and navigation, thereby fostering natural human–robot interaction.
Figure 1: Motivation of SignBot: Human-Robot Sign Language Interaction.
@@
However, no previous study has explored some methods for sign language applications. Teleoperation-based approaches [4, 5] typically rely on human manipulation, preventing robots from autonomously performing sign language. Learning-based control methods [6, 7] primarily focus on the robot’s body without addressing the complexities of dexterous hand movements. Additionally, many dexterous robotic hands have limited degrees of freedom (DoFs), and the lack of wrist flexibility further restricts the accurate expression of the rich and diverse movements required for sign language.
@@
To overcome these challenges, we introduce SignBot, an expressive robotic sign language framework designed for seamless interaction with sign language users. Figure 1 illustrates the motivation of our work. SignBot mainly consists of three key components: 1) Motion Retargeting, which maps the action sequences from human sign language datasets into a format compatible with robotic kinematics [6, 8]. 2) Policy Training (SignBot’s Cerebellum), which enables humanoid robots to first learn diverse sign language motions in a simulated environment with a decoupled policy [5, 6]. Specifically, we utilize decoupled body policies to learn the entire sign language gesture. The upper body, including the hands, learns to track the target sign language poses through imitation learning, while the lower body maintains a stable standing posture using an RL policy. 3) Sign Language Interaction (SignBot’s Cerebral) integrates a sign language translator [9], a sign language responder [10], and a sign language generator, enabling the robot to understand user expressions and respond appropriately in sign language. By combining these elements, our framework enhances real-time human-robot interaction, bridging the communication gap between sign language users and embodied robot systems.
@@
We design various experiments to verify the performance of SignBot. Experimental results show that SignBot exhibits accuracy, generalization, naturalness, and interactivity with various datasets and robots. Overall, the main contributions of our paper are as follows:
@@
    Human-Robot Interaction for Minority. We propose an interactive sign language framework that enables seamless communication between robots and the DHH community.
@@
    Precise Sign Language Execution. Our robot control policy robustly adapts to a diverse range of human sign language motions, ensuring stable and accurate execution.
@@
    Embodiment and Domain Adaptation. Signbot can be transferred to different robots, achieving sign language interaction in Sim-to-Real scenarios.
II RELATED WORK
@@
Humanoid Robotic Imitation of Human Behavior. For the imitating human behavior problem of humanoid robots, researchers often adopt a whole-body control learning paradigm [3]. This paradigm consists mainly of two approaches. One approach is to decouple the upper and lower body policies, which are responsible for managing different parts of a humanoid robot. Representative works include Exbody [6], OmniH2O [5], etc. Although upper and lower body policies are decoupled, they can still be integrated into a whole-body control paradigm. The alternative approach involves providing reference motions for the humanoid robot. Given the physical similarities to humans, a promising reference is the collection of human movements from motion datasets, such as Exbody [6], ASAP [11], H2O[12], etc. We simultaneously leverage the advantages of both methods to perform sign language. These reference motions provide rich signals for humanoid robots to imitate human-like motions.
@@
Sign Language Processing. The field encompasses two primary research directions: sign language translation (SLT), and sign language generation (SLG). SLT and SLG form complementary pathways for bidirectional communication between deaf and hearing populations, specializing in sign-to-text and text-to-sign conversion, respectively. Some studies successfully incorporated language models (LMs) pre-trained on extensive natural language corpora into SLT frameworks, yielding substantial accuracy enhancements [1]. Recent state-of-the art SLG works can be categorized into two classes: the first group of methods [13] employs diffusion models to generate sign motions conditioned on text inputs; the second group of methods [2] considers the linguistic nature of sign languages and adopts a tokenizer-LM autoregressive generation approach.
III Problem Definition
@@
Robot Learning Environment. We formulate the task of tracking human sign language motions as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $\mathcal{M}=(\mathcal{O},\mathcal{S},\mathcal{A},P_{\mathcal{T}},\mathcal{R},\mu_{0},\gamma)$, where: 1) Within the observation space $\mathcal{O}$, each observation $o_{t}\in\mathcal{O}$ consists of two components: proprioception ($o_{t}^{p}$) and goal imitation ($o_{t}^{y}$). The proprioception $o_{t}^{p}$ includes essential motion-related information such as the root state, joint positions, and joint velocities. Meanwhile, the goal imitation $o_{t}^{y}$ represents a unified encoding of the whole-body sign language pose that must be tracked during RL training. 2) $s_{t}\in\mathcal{S}$ records the complete information and environment of the robot. We summarize a state as $s_{t}=[o_{t},o_{t-1},...,o_{t-H}]$ and each $o_{t}=[o_{t}^{p},o_{t}^{y}]$. 2) $a_{t}\in\mathcal{A}$ denotes the action space, and action $a\in\mathcal{A}$ denotes the target joint positions that a PD controller uses to actuate the DoF. 3) $r_{t}=\mathcal{R}(s_{t},a_{t})$ denotes the reward functions, which typically consist of penalty, regularization, and task rewards. These reward signals determine the level of optimality in the control policy, for which we provide a detailed introduction in Appendix A.1. 4) $P_{\mathcal{T}}\in\Delta^{\mathcal{S}}_{\mathcal{S}\times\mathcal{A}}$ denotes the transition function as a mapping from state-action pairs to a distribution of future states. 5) $\mu_{0}\in\Delta^{\mathcal{S}}$ denotes the initial state distribution. 6) $\gamma\in(0,1]$ denotes the discounting factor. Under this POMDP, our goal is learning a control policy $\pi:\mathcal{S}\to\mathcal{A}$ that can maximize the discounted cumulative rewards $\sum_{t=0}^{T-1}\gamma^{t}r_{t}$.
@@
Figure 2: Overview of SignBot: The framework consists of three stages: (1) Motion Retargeting aligns human sign language gestures with the body structure of humanoid robots (Section IV-A). In addition, we use the processed mesh along with text labels to train the sign language generator model. (2) Cerebellum performs Sim2Real policy training that enables the robot to track various sign language gestures in the simulated environment and deploy the policy to real-world (Section IV-B). (3) Cerebral conducts sign language reasoning to facilitate communication with sign language users through the sign language translator, response, and generator within the cerebral (Section IV-C).
IV SignBot
In this section, we introduce the pipeline of SignBot, which is divided into three parts: 1) Motion Retargeting of the body and hands, 2) Policy Training to control the robot’s movements as "SignBot’s cerebellum", and 3) Sign Language Reasoning for comprehensive and responding users’ sign languages as "SignBot’s cerebral". Figure 2 illustrates the SignBot pipeline.
IV-A Motion Retargeting
@@
Body Retargeting. Our humanoid body retargeting system is based on the [14]. By establishing a mapping relationship between the source keypoints and the target keypoints, we follow a dual T-Pose (the standard poses of the source and target skeletons) as a spatial alignment reference. We convert the local quaternion of each joint to the form of axis angles. It is important to note that to make the robot’s sign language movements more natural, we add two additional degrees of freedom at the robot’s wrist and transform the wrists and shoulders represented by a 1D joint to a 3D joint.
@@
Hand Retargeting. To map human hand pose data to the joint positions of the Linker hand (the robot hands used by SignBot), we apply the preprocessing method from [8], adapting it specifically for the Linker hand. This process is often formulated as an optimization problem, where the difference between the keypoint vectors of the human hand model and the dexterous hand is minimized. Since the linker hand is larger than a typical dexterous hand, we adjust the scale factor in the optimization process. Additionally, we modify the regularization term to smooth the sign language movements between consecutive frames.
IV-B SignBot’s Cerebellar: Control Policy Training
@@
Within our SignBot humanoid agent, the cerebellum controls the low-level movements for performing sign language. As shown in the second stage of Figure 2, we train a control policy to enable the humanoid robot to track and imitate sign language gestures in a simulated environment. In this section, we discuss our approach from three key perspectives: observation space, decoupled policy, and reward design.
@@
Observation Space. SignBot’s observation $o_{t}\in\mathcal{O}$ consists of proprioception ($o_{t}^{p}$) and goal imitation ($o_{t}^{y}$) (Section III). Our proprioception is defined as $o_{t}^{p}\triangleq\left[\mathbf{q}_{t},\dot{\mathbf{q}_{t}},\mathbf{v}_{t},\mathbf{w}_{t},\mathbf{g}_{t},\mathbf{a}_{t-1}\right]$, which includes joint position $\mathbf{q}_{t}\in\mathbb{R}^{55}$ (DoF position), joint velocity $\dot{\mathbf{q}}_{t}\in\mathbb{R}^{55}$ (DoF velocity), root linear velocity $\mathbf{v}_{t}\in\mathbb{R}^{3}$, root angular velocity $\mathbf{w}_{t}\in\mathbb{R}^{3}$, root projected gravity $\mathbf{g}_{t}\in\mathbb{R}^{3}$, and the previous action $\mathbf{a}_{t-1}\in\mathbb{R}^{55}$. The goal observation is $\mathbf{o_{t}^{y}}\triangleq\left[\hat{\mathbf{q}}^{kp},\hat{\mathbf{q}},\dot{\hat{\mathbf{q}}}\right$], where $\hat{\mathbf{q}}^{kp}\in\mathbb{R}^{14\times 3}$ are the positions of 14 selected reference keypoints [6] to ensure that the humanoid robot and the imitation goal are oriented in the same direction, $\hat{\mathbf{q}}\in\mathbb{R}^{55\times 3}$ are the positions of all reference joints, and $\dot{\hat{\mathbf{q}}}$ is the linear velocity of the reference joints. At the same time, we generally perform domain randomization [5] in the simulation environment to ensure robustness.
@@
Decoupled Policy. Given that sign language involves precise coordination of both hand poses and full-body motion with high DoFs, learning a unified control policy is inherently challenging. More importantly, while the dual arms in the upper body can often be governed by a shared control strategy, the control approach for the lower body may vary significantly depending on the physical design of the humanoid robot. For instance, bipedal humanoid robots typically utilize RL controller, whereas wheeled robots often employ model predictive control (MPC) techniques.
@@
Motivated by recent advances in whole-body humanoid control [6], we adopt a decoupled architecture that separates the control policies of the upper and lower body.The primary objective of the upper-body policy $\pi^{upper}$ is to track the retargeted actions, while the lower-body policy $\pi^{lower}$ ensures balance in the robot’s default standing pose while adapting to the movements of $\pi^{upper}$.
@@
Where: 1) $a_{*}^{up}$ represents the retargeted action of the upper body, and $\delta$ denotes the Dirac delta function. 2) $\mathcal{D}_{f}(\cdot\|\cdot)$ refers to the $f$-divergence (e.g., KL-divergence) between two distributions. 3) The whole-body humanoid action, $a=[a^{low},a^{up}]$, is the concatenation of upper and lower body actions. 4) The reward from the humanoid control learning environment can be represented by the weighted penalty $r_{\mathfrak{P}}$, task $r_{\mathfrak{T}}$, and regularization $r_{\mathfrak{R}}$ terms: $r_{t}=\beta_{\mathfrak{T}}r_{\mathfrak{T}}+\beta_{\mathfrak{P}}r_{\mathfrak{P}}+\beta_{\mathfrak{R}}r_{\mathfrak{R}}$. In our task, task rewards ($r_{\mathfrak{T}}$) measure the robot’s performance in tracking joint motions and body velocities. Penalty rewards ($r_{\mathfrak{P}}$) serve to discourage undesirable outcomes such as falling and violating dynamic constraints like joint or torque limits. Regularization rewards ($r_{\mathfrak{R}}$) are used to align the humanoid’s sign language gestures with human preferences. Appendix A.1 specifies the reward functions.
@@
Note that to train the lower-body policy $\pi^{lower}$, we apply the PPO [15] algorithm for legged humanoid robots and MPC for wheeled humanoid robots. Appendix A.2 reports the detailed parameter settings. In this way, SignBot can be scaled to multiple embodiments, for which we demonstrate in the experiment.
IV-C SignBot’s Cerebral: Sign Language Reasoning